OpenStudy (anonymous):

I would just like assistance if that is possible. In solving the equation (x+4)(x-4)=33, Eric stated that the solution would be x+4=33=> x=29 or (x-4)=33=>x=37 However, at least one of these solutions fails to work when substituted back into the the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

jimthompson5910 (jim_thompson5910):

Eric is confusing the rule that if AB=0 then A=0 or B=0, which is known as the zero product property. Eric CANNOT say that if AB=33, then A=33 or B=33. This is simply NOT true (for every case). So what Eric needs to do is get everything to one side so that the other side is 0. Then factor and then use the zero product property to solve.

OpenStudy (anonymous):

So is the original problem x^2-16-33=0 ?

myininaya (myininaya):

very good jim

OpenStudy (anonymous):

Or do I just FOIL (x+4)(x-4) then put the 33 on the other side of that and then equal to 0?

myininaya (myininaya):

and nickie

jimthompson5910 (jim_thompson5910):

You should get x^2-16-33=0 which simplifies into x^2-49=0. Now solve this for x.

myininaya (myininaya):

(x-7)(x+7)=0 x=__ or x=____ fill in those blanks

OpenStudy (anonymous):

Thank you myininaya for the medal. Can you help me the rest of the way. oh so x=7 and x=-7

myininaya (myininaya):

np lol

jimthompson5910 (jim_thompson5910):

You are correct Nickie. The solutions are x = 7 or x = -7 Check: x = 7 (x+4)(x-4)=33 (7+4)(7-4)=33 11*3=33 33=33 Do the same with x=-7 and you'll see it works too.

myininaya (myininaya):

gj nickie :)

OpenStudy (anonymous):

Check: x=-7 (x+4)(x-4)=33 (-7+4)(-7-4)=33 -3*-11=33 33=33